OpenStudy (onepieceftw):

Help, please.. ;_; r= sqr(5)^2 + (-12)^2 r should equal sqr17 right?

OpenStudy (onepieceftw):

\[r=\sqrt{(5)^2+(-12)^2}\]

OpenStudy (freckles):

\[r=\sqrt{5^2+(-12)^2} \] well 5^2=5(5)=25 and (12)^2=12*12=144 you are to add 25 and 144 then find the square root of their sum

OpenStudy (onepieceftw):

Okay that makes sense. Thanks! The way my course made it look was really confusing.

OpenStudy (freckles):

so you know it isn't sqrt(17) right?

OpenStudy (freckles):

sqrt(169)=? ----------------- hint 13*13=169

OpenStudy (onepieceftw):

Yeah I got now. I just thought the sqr sign canceled out the ^2

OpenStudy (freckles):

nope nope remember order of operations do the operations inside the grouping symbol before taking care of the things outside the grouping symbols :) but yeah glad you got it now :)

OpenStudy (onepieceftw):

The question I was answering was: "For an angle Θ with the point (5, −12) on its terminating side, what is the value of cosine?" and the answer is 5/13

OpenStudy (freckles):

ah I knew this was related to triangles

OpenStudy (freckles):

more specifically right triangles :)

OpenStudy (freckles):

let me know if you need any further help

OpenStudy (onepieceftw):

Sure, thanks!

OpenStudy (onepieceftw):

When Θ = 5 pi over 6, what are the reference angle and the sign values for sine, cosine, and tangent?

OpenStudy (onepieceftw):

What about this one, @freckles?

OpenStudy (freckles):

the reference angle is a deg measurement between 0 and 90 or in your case the reference angle is a rad measurement between 0 and pi/2 this number can be found by finding the acute measurement between the angle given and the x-axis 5pi/6 is located: